OpenStudy (anonymous):

find the indicated outputs for f(x)= 5x2-5x ...f(0)= and f(-1) = and f(2) =

OpenStudy (anonymous):

f(0) = 10 f(-1) = 15 f(2) = 0

OpenStudy (anonymous):

thanks and this one... use the intercepts to graph this equation.. x+4y=8

OpenStudy (anonymous):

Equation is: y = -x/4 + 2 Y-intercept is 2. X-intercept is 8.

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks.. the function ,p(d) =1+ d/33 gives atmosphere at the depth in the sea (d is in feet) note thta p(0)=1 atmosphere,p(33)=2 and so on. Please find the pressure at 130 feet

OpenStudy (anonymous):

Post this as a new problem.